OpenStudy (anonymous):

A 10 gram sample of a substance that's used to detect explosives has a k-value of 0.1353. Find the subtance's half-life, in days.round to the nearest hundreds

OpenStudy (anonymous):

you shoud try to ask it in chemistry section

OpenStudy (anonymous):

ok thanks it scause i really need this answer

OpenStudy (anonymous):

do you have the formula?

OpenStudy (anonymous):

n=Noe^kt

OpenStudy (anonymous):

n is moles?

OpenStudy (anonymous):

i need to round to the nearest hundreds too

OpenStudy (anonymous):

what is the unity of k?

OpenStudy (swissgirl):

What does k represent?

OpenStudy (anonymous):

a positive constant that depends on the substance itself and on the units used to measure time

OpenStudy (anonymous):

5=10*e^(Kt), but what unity of time? seconds^-1?

OpenStudy (swissgirl):

Its in days

OpenStudy (anonymous):

there a picture of the question

OpenStudy (anonymous):

and probally the formula is e^(-kt)

OpenStudy (anonymous):

did you see the picture

OpenStudy (swissgirl):

ya its a diff question :)

OpenStudy (anonymous):

my bad ill got teh wrong one

OpenStudy (anonymous):

if k is in days^-1 just solve 5=10*e^(-kt), and you will find t

OpenStudy (anonymous):

but i dont know how to solve it either -.-

OpenStudy (anonymous):

can you help me please i really need it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):